OpenStudy (anonymous):

Medal to best answer! Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = 1/3(x - 5); (-4, 3)

OpenStudy (anonymous):

You first need to determine the slope of the given line. Rewrite that into form y = mx + b

OpenStudy (anonymous):

Then tell me the slope of the perpendicular line

OpenStudy (anonymous):

Wait so is the equation y=1/3x-1/3?

OpenStudy (anonymous):

And the slope would be -1/3.

OpenStudy (anonymous):

@JuanitaM

OpenStudy (anonymous):

the original equation is y = 1/3x -11/3 making the slope 1/3 so the perpendicular slope is -3 (remember slope is negative reciprocal)

OpenStudy (anonymous):

do you see where you made your mistake?

OpenStudy (anonymous):

But don't I have to take away the +2 from y +2 = ?

OpenStudy (anonymous):

yes which is where the -11/3 came from. i will draw it out

OpenStudy (anonymous):

OH I SEE

OpenStudy (anonymous):

Yeah my bad I'm not sure why I thought it equaled -1/3

OpenStudy (anonymous):